Calculate the pH of a buffer made from 0.28 M HNO2 and 0.23 M NO2-. The Ka for HNO2 is 4.6×10-4

1
Expert's answer
2022-05-05T04:20:03-0400

 pH = pK+ log A/HA

pKa = - logK= log 4.6•10-4 = 3.34

pH = 3.34 + log (0.23/0.28) = 3.25